# Searching from Area to Point: A Semantic Guided Framework with Geometric Consistency for Accurate Feature Matching

### v5

@@ V-D Relative Pose Estimation
1) Evaluation protocol. Following [10, 30], we report the pose estimation AUC. The camera pose is recovered by solving the essential matrix with RANSAC. Correspondences are uniformly sampled from the image, with a maximum number of 500. For ScanNet, KITTI360 and YFCC100M dataset, we report the pose AUC@$5^{\circ}/10^{\circ}/20^{\circ}$. As pose estimation is hard in MatterPort3D, we report the pose AUC@$10^{\circ}/20^{\circ}/30^{\circ}$. The comparison methods includes those in point matching experiments and other recent SOTA methods [52, 14].
TABLE VI: Relative pose estimation Results (%) on YFCC100M. Two different semantic inputs for our method are compared: SGAM using SEEM-L and SGAM using SEEM-T. The best and second results are highlighted.
Pose estimation AUC YFCC100M
@@
SP [5]+NN 16.49 30.39 45.27
SP+OANet [32] 26.82 45.04 62.17
SP+SceneFM [52] 31.45 52.27 70.31
SP+SG [30] 28.45 48.6 67.19
SP+SG+MKPC [14] 37.73+19.97% 58.03+11.02% 74.57+6.06%
ASpan [7] 38.96 59.35 75.54
SEEM-L [16]+SGAM_ASpan 39.90+2.41% 60.36+1.70% 76.34+1.06%
SEEM-T [16]+SGAM_ASpan 39.77+2.08% 60.24+1.50% 76.21+0.89%
@@
COTR [37] 39.73 59.12 74.27
SEEM-T+SGAM_COTR 39.94+0.53% 60.89+2.99% 74.65+0.51%
SEEM-L+SGAM_COTR 40.19+1.16% 60.31+1.97% 74.76+0.66%
@@
2) Indoor Results. The pose AUC results in indoor scene are summarised in Tab. V-B and Tab. V. It is evident that our method significantly improves performance across all baselines, attributable to both enhanced matching accuracy and the even matching distribution. Particularly noteworthy is the impressive precision improvement achieved on the challenging MatterPort3D dataset, underscoring the superiority of our method in tackling difficult matching scenarios. Furthermore, similar to the point matching experiment, higher semantic precision leads to improved geometry estimation. The recent SEEM is able to offer enough accurate semantic prior for our method, even it is trained with COCO [48] labels. Additionally, in Table V, we compare our method with other leading matching approaches on the standard ScanNet-1500 benchmark. Since this benchmark lacks semantic labels, we employ SEEM-L to perform semantic segmentation for our method. In this experiment, we further combine SGAM with a dense matching method, DKM [40], whose performance is also significantly boosted by our approach. The results demonstrate that our method outperforms other approaches by a considerable margin, achieving the highest accuracy and highlighting its effectiveness in practical applications. In particular, there is another two-stage matching method similar to our method, MKPC [14], which achieves co-visible area to match by stage-one point matching. However, its accuracy gains for the baseline is very limited. This limitation may be attributed to the fact that MKPC can only achieve single area match for each pair of images. Moreover, its area matching accuracy is dependent on precise point matching, which is challenging to achieve in complex indoor scenes. In contrast, SGAM is particularly well-suited for indoor scenes due to their abundance of semantic information, making it more effective in this context.
@@
3) Outdoor Results. The results on KITTI360 dataset are reported in Tab. V-C. Our method greatly enhances the precision of pose estimation for all baselines, affirming the effectiveness and robustness of our method in outdoor driving scenes. It is important to highlight that the impact of different semantic inputs on performance is minimal here. This is primarily due to the reduced semantic complexity of driving scenes compared to indoor scenes, wherein the SEEM-L backbone can yield accurate semantic segmentation outcomes. Thus, SGAM demonstrates the capability to achieve precise pose estimation when SEEM is employed to obtain semantic prior, thereby showcasing the practicality of our approach.
The Tab. VI reports the results on YFCC100M dataset. Similar to indoor experiments, we obtain semantic input using SEEM with two backbones to study the impact of different semantic inputs on performance. As we can seen in this table, our SGAM is able to boost the performance for all baselines. However, the improvements in this outdoor scene are limited (up to $+2.41\%$), compared to indoor and driving scenes. The underlying reason is that the semantic information in this scene is very limited, leading to semantic segmentation at large granularity, e.g. segmentation only contains labels like ‘sky’, ‘building’ and ‘people’. Thus, SGAM attains few area matches, usually covering almost the whole image, and therefore has limited advantages for subsequent point matching. In contrast, MKPC is more suitable to this dataset and achieves more improvement. However, SGAM does outperform MKPC in accuracy, combined with all baselines. Additionally, we provide some qualitative comparison examples in Fig. 5, where the accuracy of our SGAM against the baselines is improved.
@@
TABLE VII: Area matching performance on ScanNet. The area matching results (%) of SAM and SGAM combined with different point matchers under three matching difficulties and three semantic input settings in ScanNet are reported. The best and second results under each semantic input setting are highlighted.
@@ V-E Area Matching
2) Results. The area matching results are summarised in Tab. VII. The threshold $t$ of AMP is set as $0.7$. We analyze the outcomes of SAM and SGAM when combined with SOTA detector-free matchers. Within the table, the precision of area matching in SAM decreases as the matching difficulty increases. When the semantic input is accurate (GT), the AMP values demonstrate that most areas are accurately matched under all conditions. However, as semantic precision decreases, our method also experiences a decrease in precision, which is more pronounced in large FD. This demonstrates the main limitation of our method, i.e. the heavy reliance on semantic, which is discussed in detail in Sec. V-K. Notably, the utilization of SGAM enhances the accuracy of area matching in all scenarios, highlighting the importance of GAM in area matching. Different point matchers also result in different regional matching precision, but the overall difference is small, proving the compatibility of our method for point matchers.
V-F Understanding SAM
SAM contains matching two kinds of areas: the semantic object area (SOA) and the semantic intersection area (SIA). To assess the importance of these two areas on area matching, we designed experiments to evaluate their quantities and matching accuracy in ScanNet image pairs with FD@5/10/30. To further investigate the performance under different semantic input accuracy, SAM takes three semantic segmentation as input, including ground truth (GT sem.) and segmentation by SEEM with two backbones (SEEM-L/T Sem.). The results are summarised in the Tab. VIII, including the SOA Match and SIA Match under different FD and semantic input. It can be seen that SOA matches are more accurate and robust against semantic precision compared to SIA. This can be attributed to the better stability of the centered object semantic against various matching noises. On the other hand, the precision of SAM is limited by semantic precision. As shown in the table, the accuracy of area matching decreases with semantic accuracy, and the decrease is more significant at large FD. This limitation is discussed in detail in Sec. V-K. In addition, both two areas are frequently involved in the matching and sufficient number of area matches is also important for downstream tasks, indicating their importance for SGAM.
Figure 7: Visualization of GP. The cases processed by GP, which can predict the true matches (boxes of the same color) from the doubtful candidates (yellow boxes) in semantic ambiguity.
TABLE VIII: Area matching performance on ScanNet of two SAM areas and GP. We construct area matching experiments for matching of two semantic areas, along with GP integrated with four point matchers. The effect of three different semantic inputs on performance is tested. AOR and AMP (with threshold $t=0.7$) under different matching difficulties (each with 1500 image pairs) are reported along with the area number per image (Num). The best and second results under each semantic input setting and FD setting are highlighted.
@@ V-F Understanding SAM
SEEM-T Sem. SOA Match 80.97 81.31 3.78 75.12 71.21 3.37 63.73 47.10 2.69
SIA Match 70.57 63.49 1.82 64.13 52.23 1.63 53.70 38.45 1.47
GP_ASpan 81.39 83.19 0.77 77.17 73.19 1.15 64.89 48.93 1.60
GP_QuadT 81.09 82.79 77.59 72.38 64.36 48.76
GP_LoFTR 81.74 82.91 77.45 72.12 64.29 48.59
GP_COTR 81.59 82.63 76.93 71.61 64.17 48.15
Figure 8: Visualization of GR. After GR, many false and inaccurate area matches (boxes of the same color) are rejected. Only reliable area matches (highlighted by yellow boxes) are left, leading to high matching accuracy for point matching.
V-G Understanding GAM
1) GP Precision. This section focuses on examining the area matching performance of GP on ScanNet [43] across three difficulty levels and semantic input (GT and SEEM-T/L), each consisting of 1500 image pairs. The results are shown in Tab. VIII. It can be seen that the area matching precision of GP surpasses that of SAM under different matching difficulties. Similar to SAM, the precision of GP decreases with inaccurate semantic input, but it can establish more accurate area matches than SAM in all semantic input cases. These observations confirm the effectiveness of GP. Furthermore, the choice of point matchers influences the performance of GP, with improved point matching leading to increased accuracy in area matching. Notably, the doubtful area count per image indicates non-trivial semantic ambiguity within SAM, which becomes more prevalent with increasing matching difficulty. In conclusion, GP plays a crucial role in SGAM by addressing semantic ambiguity and enhancing area matching performance. The visualization of GP are shown in Fig. 7.
2) Ablation Study of GR Parameter. In order to thoroughly examine the influence of the parameter $\phi$ in GR on pose estimation performance, we conduct experiments on ScanNet. We evaluate two difficulty levels, FD@10 and FD@30, each consisting of 1500 image pairs. The obtained results are depicted in Fig. 9. It is evident that the choice of $\phi$ has minimal impact when the frame difference is 10. In scenes where the matching difficulty is not too high, an adequate number of area matches are identified. With a smaller $\phi$ value, accurate area matches are selected, (thanks to our GMC module which ensures the even distribution of matches) leading to improved performance. However, as the difficulty of area matching increases under FD@30, a smaller $\phi$ value can cause point matches to be spatially concentrated, resulting in planar degeneration in certain cases. Hence, selecting an appropriate $\phi$ is crucial in challenging matching scenarios. Based on empirical findings, we establish $\phi=0.5$ as the default value. We visualize the GR in Fig. 8.
@@ V-G Understanding GAM
TABLE IX: Comparison experiment on maximum correspondence number (MCN). The pose estimation results on ScanNet with FD@10 and YFCC100M are reported. Our method is combined with ASpan and takes semantic input from SEEM-T. We show the improvement of our method on the accuracy of ASpan.
MCN Method ScanNet-FD@10 YFCC100M
AUC@5$\uparrow$ AUC@10$\uparrow$ AUC@20$\uparrow$ AUC@5$\uparrow$ AUC@10$\uparrow$ AUC@20$\uparrow$
500 ASpan 58.51 70.42 79.84 38.96 59.35 75.54
SGAM_ASpan 60.78+3.88% 74.24+5.42% 84.53+5.87% 39.77+2.08% 60.24+1.50% 76.21+0.89%
800 ASpan 58.02 68.45 78.41 39.35 59.72 75.67
SGAM_ASpan 60.56+4.38% 71.17+3.97% 79.84+1.82% 40.05+1.78% 60.53+1.36% 76.54+1.15%
1000 ASpan 57.81 67.66 77.59 39.18 59.54 75.60
SGAM_ASpan 60.24+4.20% 72.65+7.38% 80.77+4.10% 40.55+3.50% 60.84+2.18% 76.78+1.56%
V-H Understanding Global Match Collection
@@
Global Match Collection (GMC) is another important module of our method, which ensures widely distributed point matches, particularly in less semantic scenes. To demonstrate the contribution of this module in our method, we conducted an ablation study on the ScanNet-1500 benchmark. The results are summarised in Fig. 10. The main parameter of our study is the size proportion threshold ($T_{SP}$), which represents the proportion of the image occupied by matched areas in the image pair. When $T_{SP}=0$, no GMC is performed, and when $T_{SP}=1$, all image pairs adopt the GMC module. First, we demonstrate the relationship between $T_{SP}$ and the collection proportion, which represents the percentage of cases that adopt the GMC among all the pairs (Fig. 10(a)). The relationship is close to linear, which is consistent with common sense. Next, we display the pose estimation performance (AUC@$5^{\circ}/10^{\circ}$) under different size proportion thresholds (Fig. 10(b)). In the figure, the performance increases significantly as the threshold changes from $0.1$ to $0.3$. This is because the GMC module can significantly improve the distribution of matches, especially when there are only a few area matches established. It is worth noting that even without GMC, our method can slightly improve the precision of the baseline (AUC@$5^{\circ}$: 25.89 vs. 25.78). However, the GMC module brings better performance. As the threshold further increases, the performance improvement levels off because the area matches already cover most of the overlap in the image pair. Therefore, to achieve better performance while considering the computation cost of the GMC module (equal to one area match with images resized to the default size), the size proportion threshold can be set to 0.6. Moreover, we visualize the affects of GMC in Fig. 10 (c), where the distribution of matches is more uniform in the overlap area after applying GMC. Another concern about GMS is that it achieves more but may be duplicate correspondences, due to the overlaps between area matches and GMS. However, as we uniformly sample no more than the Maximum Correspondence Number (MCN) of points in the image space for pose estimation, duplicate correspondences are removed. Additionally, we conduct experiments on both ScanNet (FD@10, 1.5k image pairs) and YFCC100M, to investigate the influence of the MCN on the pose estimation, which is set as 500, 800 and 1000 respectively. The results are reported in Table IX. It is evident that the MCN have slight impact on pose estimation and SGAM demonstrates improvement across all settings.
@@
Figure 10: Ablation study of global match collection. We conduct the ablation study to evaluate the effectiveness of global match collection module on the ScanNet-1500 benchmark. (a) The relationship between the size proportion threshold and the number of cases using the global match collection module (collection proportion). (b) The relationship between the size proportion threshold and the pose estimation precision (Pose AUC@5/10). (c) The qualitative cases before and after global match collection. The red dash boxes indicate the area matches.
TABLE X: Decomposing component ablation study. We conduct the decomposing component experiment of SGAM_ASpan on ScanNet with FD@15, using 1k image pairs. The area matching and pose estimation performance are reported. The numbers of area matches (Num) are also reported.
SOA SIA GP GR GMC AOR $\uparrow$ AMP $\uparrow$ Num AUC@5 $\uparrow$ AUC@10 $\uparrow$ AUC@20 $\uparrow$
79.39 78.88 2.74 32.96 43.77 53.80
74.64 69.42 2.11 31.72 40.66 46.89
77.41 74.95 4.53 34.02 48.26 52.94
79.03 76.30 5.57 39.54 52.97 64.54
79.01 78.00 3.73 49.24 62.01 73.32
79.18 78.27 4.01 49.73 62.58 73.17
79.18 78.27 4.01 50.50 63.38 74.75
V-I Decomposing Ablation Study
We also conducted a dedicated experiment to systematically decompose the components of our approach, evaluating the area matching and pose estimation performance. The threshold of AMP is set as $t=0.7$. We sample 1500 image pair with frame difference is 15 (FD@15) from ScanNet for this experiment. Our SGAM is combined with ASpan. The results are reported in Tab. X. It can be seen that both area matching and pose estimation accuracy improve as the completeness of our method increases. This finding confirms the effectiveness of all the components within our SGAM approach.
TABLE XI: Time Consumption Comparison. The experiment is conducted on ScanNet with FD@10. The time consumption of each component of our method with specific input size is reported. Different time consumption comes from different baselines coupled with our method are investigated as well. The time of baselines are also reported.
Time/s Input Size
640$\times$480 256$\times$256 640$\times$480
PM SAM GP GR GMC SGAM PM
ASpan 0.62 0.042 0.20 0.021 0.88 0.19
QuadT 0.040 0.17 0.023 0.85 0.18
LoFTR 0.041 0.19 0.018 0.86 0.18
COTR 2.54 23.85 2.13 29.14 56.04
V-J Discussion on the Computational Complexity
In this section, we discuss in detail the theoretical computational complexity of each component of our proposed approach, including SAM, GP and GR. Moreover, we conduct an experiment to count the average running time per image pair of each component of our method in practice. We collect 1500 sets of image pairs from ScanNet with FD@10 for this experiment. Four baseline point matchers are combined for time comparison. The results are reported in Tab. XI. This experiment is run on a Intel Xeon Silver 4314 CPU and a GeForce RTX 4090 GPU.
1) Computational Complexity of SAM. SAM mainly includes area detection and description for two semantic areas (SOA and SIA). For SOA, the size of the algorithm is related to the number of semantic categories in the image pair ($N_{sem}$). Thus the computational complexity for this part is $O(N_{sem})$. The SIA part involves a sliding window algorithm on image. Its computational complexity is $O((W_{I}-W_{w})*(H_{I}-H_{w})/s^{2})$, where $W_{I},H_{I}$ are the Width and Height of the Image, $W_{w},H_{w}$ are Width and Height of window and $s$ is the sliding step. As the window size is large (please refer to Sec. IV-A2), the time consumption of this part is acceptable. 2) Computational Complexity of GP. Given $H$ doubtful areas in $I_{0}$ and $R$ in $I_{1}$, the GP can determine area matches through point matching within areas. Relying on correspondences, however, leads to multiple times point matching for all area match possibilities ($H\times R$ times) in the single image pair. At the same time, the calculation of $P({As}_{l})$ in Eq. 9 also need to be performed $L=\frac{H!}{(H-R)!}$ times. Thus, for $P({As_{l}})$ calculation, its computational complexity varies from $O(N)$ (when $R=1$) to $O(N!)$ (when $R=H-1$), depending on the area number. However, as we only perform this prediction when semantic ambiguity occurs, the practical time cost is acceptable, as shown in Tab. XI.
3) Computational Complexity of GR. In GR, point matching inside area matches are performed to compute geometry consistency. This inside-area matching is the key of our A2PM framework, which is equivalent to decomposing a large matching problem (the full-image matching) into multiple smaller matching problems. The computational complexity of widely-used vanilla Transformer in SOTA matching method [37] is $O(N^{2})$, where $N$ is the input size. Thus, this decomposing of matching in A2PM is more efficient than direct matching, when the area size is smaller enough than the original image size. Specifically, suppose the image size is $W_{I}\times H_{I}$, area size is $W_{a}\times H_{a}$ and area match number is $N_{a}$. Then the computational complexity of original point matcher is $O((W_{I}\times H_{I})^{2})$, while the computational complexity of A2PM is $O(N_{a}\times(W_{a}\times H_{a})^{2})$. Thus, our A2PM is more effective than the original point matcher, when $(W_{I}\times H_{I})^{2}/(W_{a}\times H_{a})^{2}\geq N_{a}$. Some recent methods use linear Transformer whose computational complexity is $O(N)$. In this case, our A2PM is more effective than the original point matcher, when $(W_{I}\times H_{I})/(W_{a}\times H_{a})\geq N_{a}$. The specific time consumption of our method is reported in Tab. XI for ScanNet. The area size is $256\times 256$ and the image size is $640\times 480$ in the experiments. It can be seen that our GR costs similar time as the original PM, when linear Transformer is used (ASpan, QuadT and LoFTR). However, the overall time consumption of our SGAM is more than direct PM. This is primarily attributed to the slower speed of SAM, as we have not yet optimized it for efficiency in engineering. On the other hand, our method is faster than direct PM with vanilla Transformer (COTR), because of its second-order computational complexity.
Figure 11: Qualitative Comparison on challenging scenes. The visual comparison between our method and ASpan in challenging scenes. The wrong and correct matches under the same threshold are labeled respectively.
V-K Advantages and Limitations
As semantic possesses consistency against various matching noises, e.g. illumination, viewpoint and scale changes between images, our SGAM is able to find accurate area matches in challenging scenes. Then, the precision of inside-area point matching is significantly boosted, due to noise removal and higher resolution of these areas, which is the main advantage of our method. We also provide qualitative comparison results for five hard scenarios in Fig. 11 to demonstrate the superiority of SGAM. However, heavy reliance on semantics also results in limitations of SGAM. First, the semantic segmentation accuracy impacts the performance of our method. Especially for area matching, our SAM and GP exhibits non-negligible decreases in precision (Tab. VII and Tab. VIII). This is because the area detection and description in SAM both assume accurate semantic input. Thus the under-splitting, over-splitting, and multiple semantics in inferior semantic segmentation leads to reduced performance of SAM and large doubtful area numbers for GP. However, it is noteworthy that SGAM is still able to improve the point matching and pose estimation performance with SEEM-L/T in our experiments. This highlights the fact that the advantages of SGAM are still significant in the presence of less accurate semantic inputs and implies the potential of our A2PM framework. The second limitation of SGAM is related to the spatial granularity of semantic categories. For example, when a single semantic entity dominates the image, it is difficult for SAM to find areas with clustered features. Hence the effectiveness of SGAM is restricted in some scenes, such as YFCC100M. However, in such scenes, the A2PM framework still benefits feature matching, but the area matches need to be established by other approaches, like overlap estimation [53, 14]. In our future work, we will focus on area matching without semantic prior, which may work well in more general application scenes.
VI Conclusion
From a unified searching perspective, this paper proposes semantic area matches, which are matched areas with prominent semantic, to serve as the proper search space for precise feature matching. This search space effectively reduces the matching difficulty and provides point matchers with high resolution input, ultimately improving the matching performance. We propose the A2PM framework as the corresponding search strategy for this search space, which hierarchically divides the feature matching into two search phases: establishing semantic area matches on the entire image and finding point matches from these area matches. To implement the A2PM framework, we propose SGAM comprises SAM and GAM, to leverage both semantic information and geometry constraint of images. The SAM performs rough area matching based on semantic cues. On the ther hand, the GAM, integrated with an off-the-shelf point matcher, achieves accurate area and point matches based on the geometry consistency. Extensive experiments demonstrate the effectiveness of our approaches, which significantly outperform other SOTA methods in both image matching task and down-stream pose estimation task.

### v6

@@ IV-C Relative Pose Estimation
1) Evaluation protocol. Following [9], we report the pose estimation AUC, which reveals the proportion of correct pose estimation among all estimations. The camera pose is recovered by solving the essential matrix with RANSAC. Correspondences are uniformly sampled from the image, with a maximum number of 500. SGAM is also combined with three kinds of point matchers to demonstrate the advantages adopting the proposed search space. We replace COTR with DKM [12] as the proxy of dense methods, due to its impressive performance. For ScanNet, KITTI360 and YFCC100M dataset, we report the pose AUC@$5^{\circ}/10^{\circ}/20^{\circ}$. As pose estimation is hard in MatterPort3D, we report the pose AUC@$10^{\circ}/20^{\circ}/30^{\circ}$.
TABLE V: Relative pose estimation Results (%) on YFCC100M. Two different semantic inputs for our method are compared: SGAM using SEEM-L and SGAM using SEEM-T.
Pose Estimation YFCC100M
@@
PATS[15] 39.25 60.77 76.38
Sparse SP+OANet [31] 26.82 45.04 62.17
SP+SG [29] 28.45 48.60 67.19
OETR [14]+SP+SG 31.51+10.76% 50.61+4.14% 70.02+4.21%
SEEM-L [18]+SGAM_SP+SG 29.54+3.83% 50.48+3.87% 69.64+3.65%
SEEM-T [18]+SGAM_SP+SG 29.14+2.43% 50.01+2.90% 68.26+1.59%
Semi-Dense ASpan [10] 38.96 59.35 75.54
OETR+ASpan 39.31+0.90% 60.13+1.31% 76.22+0.90%
SEEM-L+SGAM_ASpan 39.90+2.41% 60.36+1.70% 76.34+1.06%
SEEM-T+SGAM_ASpan 39.77+2.08% 60.24+1.50% 76.21+0.89%
@@
Dense DKM [12] 43.12 63.78 79.13
OETR+DKM 43.28+0.37% 64.27+0.77% 79.34+0.27%
SEEM-L+SGAM_DKM 43.77+1.51% 64.12+0.53% 79.94+1.01%
SEEM-T+SGAM_DKM 43.56+1.02% 63.99+0.33% 79.77+0.81%
@@
2) Indoor Results. The pose AUC results in indoor scene are summarised in Tab. II and Tab. IV. In Tab. II, it can be seen that our method consistently improves performance across all point matching baselines, indicating the versatility of our approach. The impressive precision improvement achieved on the challenging MatterPort3D dataset underscores the ability of our method in tackling difficult matching scenarios. Furthermore, similar to the point matching experiment, higher semantic precision leads to improved geometry estimation. The recent SEEM is able to offer enough accurate semantic prior for our method, even hand-crafted feature-based method is applied in SGAM. Additionally, in Table IV, we compare SGAM with other leading approaches on ScanNet1500 benchmark. The results demonstrate that our method boosts all the baselines by a considerable margin, achieving the highest accuracy. Meanwhile, the co-visible area estimation method, MKPC [13], achieves very limited improvement for the sparse baseline, compared to ours (e.g. $16.18_{-0.12\%}$ vs. $17.33_{+6.98\%}$). This may be interpreted as MKPC can only achieve single area match for each pair of images. Moreover, its area matching accuracy is dependent on precise point matching, which is challenging to achieve in complex indoor scenes. In contrast, SGAM is particularly well-suited for indoor scenes due to the abundance of semantic information, making it more effective to reduce redundant computation and improve the accuracy.
@@
3) Outdoor Results. The results on KITTI360 dataset are reported in Tab. III. Our method greatly enhances the precision of pose estimation for all baselines, affirming the effectiveness and robustness of our method in outdoor driving scenes. It is important to highlight that the impact of different semantic inputs on performance is minimal here. This is primarily due to the reduced semantic complexity of driving scenes compared to indoor scenes, wherein the SEEM-L backbone can yield accurate semantic segmentation outcomes. The table also includes a comparison with SFD2 [16], which incorporates semantic perception trained specifically for driving scenes. While the integration of semantics in SFD2 benefits matching in difficult scenes [16], enhancing features of detailed search spaces by semantic remains a challenge. Conversely, SGAM exhibits superior performance, underscoring the effectiveness of our semantic-friendly search space.
The Tab. V reports the results on YFCC100M dataset. As we can seen in the table, SGAM is able to boost the performance for all baselines, including sparse, semi-dense and dense methods. However, the improvements in this outdoor scene are somewhat restricted (up to $+2.41\%$), compared to indoor and driving scenes. This limitation arises from the scarcity of semantic information in the scene, leading to semantic segmentation at a broad granularity level, e.g. segmentation only contains labels like ‘sky’, ‘building’ and ‘people’. Consequently, SGAM generates few area matches, typically encompassing almost the entire image. Conversely, the co-visible matching method, OETR, can match co-visible areas without considering semantics. Therefore, OETR can reduce the redundant computation more effectively and surpass SGAM in terms of accuracy for sparse baseline. Despite that, SGAM demonstrates a comparable improvement against OETR for semi-dense and dense baselines, suggesting that the high accuracy of baselines may compensate for shortcomings in area matching accuracy. Additionally, we provide some qualitative comparison examples in Fig. 4.
@@
TABLE VI: Area matching performance on ScanNet. The area matching results (%) of SAM and SGAM combined with different point matchers under three matching difficulties and three semantic input settings in ScanNet are reported. The best and second results under each semantic input setting are highlighted.
@@ IV-D Area Matching
2) Results. The area matching results are summarised in Tab. VI. The threshold $t$ of AMP is set as $0.7$. We analyze the outcomes of SAM and SGAM when combined with SOTA detector-free matchers. Within the table, the precision of area matching in SAM decreases as the matching difficulty increases. When the semantic input is accurate (GT), the AMP values demonstrate that most areas are accurately matched under all conditions. However, as semantic precision decreases, our method also experiences a decrease in precision, which is more pronounced in large FD. This demonstrates the main limitation of our method, i.e. the heavy reliance on semantic, which is discussed in detail in Sec. VI-E. Notably, the utilization of SGAM enhances the accuracy of area matching in all scenarios, highlighting the importance of GAM in area matching. Different point matchers also result in different regional matching precision, but the overall difference is small, proving the compatibility of our method for point matchers.
IV-E Understanding SAM
SAM contains matching two kinds of areas: the semantic object area (SOA) and the semantic intersection area (SIA). To assess the importance of these two areas on area matching, we designed experiments to evaluate their quantities and matching accuracy in ScanNet image pairs with FD@5/10/30. To further investigate the performance under different semantic input accuracy, SAM takes two semantic segmentation as input, including ground truth (GT sem.) and segmentation by SEEM with large backbones (SEEM-L Sem.). The results are summarised in the Tab. VII, including the SOA Match and SIA Match under different FD and semantic input. It can be seen that SOA matches are more accurate and robust against semantic precision compared to SIA. This can be attributed to the better stability of the centered object semantic against various matching noises. On the other hand, the precision of SAM is limited by semantic precision. As shown in the table, the accuracy of area matching decreases with semantic accuracy, and the decrease is more significant at large FD. This limitation is discussed in detail in Sec. VI-E. In addition, both two areas are frequently involved in matching and sufficient number of area matches is also important for downstream tasks, indicating their importance for SGAM.
TABLE VII: Area matching performance of two SAM areas and GP. We construct area matching experiments on ScanNet for matching of two semantic areas, along with GP integrated with four point matchers. The effect of two different semantic inputs is also evaluated. AOR and AMP (with threshold $t=0.7$) under different matching difficulties (each with 1500 image pairs) are reported along with the area number per image (Num). The best and second results under each semantic input setting and FD setting are highlighted.
@@ IV-E Understanding SAM
IV-F Understanding GAM
1) GP Precision. This section focuses on examining the area matching performance of GP on ScanNet [42] across three difficulty levels and two semantic input (GT and SEEM-L), each consisting of 1500 image pairs. The results are shown in Tab. VII. It can be seen that the area matching precision of GP surpasses that of SAM under different matching difficulties. Similar to SAM, the precision of GP decreases with inaccurate semantic input, but it can establish more accurate area matches than SAM in all semantic input cases. These observations confirm the effectiveness of GP. Furthermore, the choice of point matchers influences the performance of GP, with improved point matching leading to increased accuracy in area matching. Notably, the doubtful area count per image indicates non-trivial semantic ambiguity within SAM, which becomes more prevalent with increasing matching difficulty. In conclusion, GP plays a crucial role in SGAM by addressing semantic ambiguity and enhancing area matching performance. The visualization of GP are shown in Fig. 10 of the appendix.
2) Ablation Study of GR Parameter. In order to thoroughly examine the influence of the parameter $\phi$ in GR on pose estimation performance, we conduct experiments on ScanNet. We evaluate two difficulty levels, FD@10 and FD@30, each consisting of 1500 image pairs. The obtained results are depicted in Fig. 6. It is evident that the choice of $\phi$ has minimal impact when the frame difference is 10. In scenes where the matching difficulty is not too high, an adequate number of area matches are identified. With a smaller $\phi$ value, accurate area matches are selected, (thanks to our GMC module which ensures the even distribution of matches) leading to improved performance. However, as the difficulty of area matching increases under FD@30, a smaller $\phi$ value can cause point matches to be spatially concentrated, resulting in planar degeneration in certain cases. Hence, selecting an appropriate $\phi$ is crucial in challenging matching scenarios. Based on empirical findings, we establish $\phi=0.5$ as the default value. We visualize the GR in Fig. 9 of the appendix.
@@ IV-F Understanding GAM
IV-G Understanding Global Match Collection
@@
Global Match Collection (GMC) is another important module of our method, which ensures widely distributed point matches, particularly in less semantic scenes. To demonstrate the contribution of this module in our method, we conducted an ablation study on the ScanNet1500 benchmark. The results are summarised in Fig. 7. The main parameter of our study is the size proportion threshold ($T_{SP}$), which represents the proportion of the image occupied by matched areas in the image pair. When $T_{SP}=0$, no GMC is performed, and when $T_{SP}=1$, all image pairs adopt the GMC module. First, we demonstrate the relationship between $T_{SP}$ and the collection proportion, which represents the percentage of cases that adopt the GMC among all the pairs (Fig. 7(a)). The collection proportion increases fast when size proportion threshold is small, indicating few-area-match pairs benefit from GMC. Next, we display the pose estimation performance (AUC@$5^{\circ}/10^{\circ}$) under different size proportion thresholds (Fig. 7(b)). In the figure, the performance increases significantly as the threshold changes from $0.1$ to $0.3$. This is because the GMC module can significantly improve the distribution of matches, especially when there are only a few area matches established. It is worth noting that even without GMC, our method can slightly improve the precision of the baseline (AUC@$5^{\circ}$: 25.89 vs. 25.78). However, the GMC module brings better performance. As the threshold further increases, the performance improvement levels off because the area matches already cover most of the overlap in the image pair. Therefore, to achieve better performance while considering the computation cost of the GMC module (equal to one area match with images resized to the default size), the size proportion threshold can be set to 0.6. Moreover, we visualize the affects of GMC in Fig. 7 (c), where the distribution of matches is more uniform in the overlap area after applying GMC.
@@
Figure 7: Ablation study of global match collection. We conduct the ablation study to evaluate the effectiveness of global match collection module on the ScanNet1500 benchmark. (a) The relationship between the size proportion threshold and the number of cases using the global match collection module (collection proportion). (b) The relationship between the size proportion threshold and the pose estimation precision (Pose AUC@5/10). (c) The qualitative cases before and after global match collection. The red dash boxes indicate the area matches.
IV-H Running Time Comparison
The proposed A2PM framework inherently decomposes the original matching task into multiple simpler matching tasks, leading to an inevitable increase in the time cost of SGAM. To demonstrate this, we conducted experiments on the YFCC100M dataset to compare the specific time costs among our method, the original point matching method, and another two-stage matching method, OETR. These experiments were performed on an Intel Xeon Silver 4314 CPU and a GeForce RTX 4090 GPU, and the results are presented in Table VIII. It is evident from the results that SGAM amplifies the time cost across all baselines, given that point matching is executed multiple times within the A2PM framework. However, SGAM demonstrates a similar time cost to OETR, although OETR conducts point matching only once. This can be attributed to the semantic-aware search space utilized by SGAM, which results in a lightweight hand-crafted approach, contrasting with the computationally intensive nature of the learning approach in OETR. Exploring the potential for parallel execution of multiple point matching in A2PM, akin to PATS (multiple patch matching operations are accelerated by CUDA), has the promise of substantially reducing the overall computational time. This avenue is one that we intend to investigate in our future research. The detailed analyze of theoretical computational complexity of each part of SGAM is left in Sec. VI-D of the appendix, along with the corresponding time costs.
TABLE VIII: Running time comparison (s). We construct experiments on YFCC100M to compare the time consuming between original point matchers and SGAM intergrated with them, along with another two-stage matching method.
SP+SG ASpan QT LoFTR COTR PATS
time of original 0.11 0.36 0.37 0.34 7.64 0.94
time w/ SGAM 1.04 1.42 1.49 1.37 22.52 -
time w/ OETR 0.84 1.03 1.14 0.95 8.32 -
V Conclusion
To better incorporate semantic robustness into the coarse-to-fine feature matching, this study proposes semantic area matches as an intermediate search space for precise feature matching. The search space represents areas within images that exhibit prominent semantic features. By utilizing this search space, redundant computations are reduced, and the following point matcher receives high-resolution input, thereby improving overall matching performance. Aligned with the search space, the A2PM framework is introduced to hierarchically divide feature matching into two phases: first, establishing semantic area matches across the images, and then finding point matches within these area pairs. To implement the A2PM framework, we further propose SGAM method, comprising SAM and GAM, which leverages both semantic information and geometric constraints within the images. SAM conducts putative area matching based on large language model-powered semantic perception, while GAM, in conjunction with a point matcher, achieves precise area and point matches by ensuring geometric consistency. Extensive experiments validate the effectiveness of our approach, enhancing performance across sparse, semi-dense, and dense matching methods in point matching and downstream pose estimation tasks.